# Geometric Criticality in Scale-Invariant Networks

### v1

@@ Abstract
Dimension in physical systems determines universal properties at criticality. Yet, the impact of structural perturbations on dimensionality remains largely unexplored. Here, we characterize the attraction basins of structural fixed points in scale-invariant networks from a renormalization group perspective, demonstrating that basin stability connects to a structural phase transition. This topology-dependent effect, which we term geometric criticality, triggers a geometric breakdown hereto unknown, which induces non-trivial fractal dimensions and unveils hidden flows toward unstable structural fixed points. Our systematic study of how networks and lattices respond to disorder paves the way for future analysis of non-ergodic behavior induced by quenched disorder.
@@
Topological shortcuts and structural sparsity shape complex dynamics across a broad spectrum of physical systems Betzel and Bassett (2018); Martin et al. (2015); Simonov and Goodwin (2020). Shortcuts are key, for instance, in shaping brain functionality Sporns (2010) or in improving the speed, efficiency, robustness, and functional diversity across neural networks Kaiser and Hilgetag (2006); Bullmore and Sporns (2009); Betzel and Bassett (2017). Similar concepts pervade other areas, such as wormholes in general relativity Schein and Aichelburg (1996) or wormhole attacks in wireless sensor networks Chen et al. (2014). In contrast, sparsity – for example, selective bond breaking in zeolites Morris and Čejka (2015); Roth and et al. (2013) – limits connectivity and tends to suppress large-scale organization. Vacancies – missing links in real materials – often compromise emergent phases by disrupting connectivity and collective behavior Simonov and Goodwin (2020). Such defects are known to modify electronic, magnetic, and structural properties in oxides, alloys, and skyrmion crystals Chen (2016); Sun (2021); Zhou (2023). Despite their relevance, understanding whether shortcuts and sparsity lead to structural transitions in complex systems remains a crucial open challenge.
@@
In the context of complex networks, shortcuts were first formalized in the celebrated Watts-Strogatz (WS) model Watts and Strogatz (1998). There, a simple ring lattice is ‘rewired’ to introduce increasing amounts of disorder, allowing for a middle ground between regular and random networks while preserving high clustering and a very slowly growing diameter Newman and Watts (1999); Newman (2000). Extensions of the WS model, where the rewiring probability decays with distance, have further demonstrated that such shortcuts can modify the spectral dimension of the underlying graph Millán et al. (2021). However, do small-world effects merely induce smooth crossovers towards randomness or possibly lead to genuine structural phase transitions?
@@
Dilute lattice models have long suggested that structural perturbations may influence universality classes Vojta (2006). Griffiths theoretically demonstrated that disordered systems could exhibit non-analytic free energy over an extended temperature range Griffiths (1969). Understanding Griffiths phases is essential for comprehending many other disordered systems, such as dilute ferroelectric materials Westphal et al. (1992); Timonin (1997). However, no general framework has emerged to predict when and how such quenched disorder induces non-ergodic behavior in arbitrary network architectures.
@@
This motivated recent work on intrinsically scale-invariant networks, where structural fixed points were defined through the Laplacian Renormalization Group (LRG) Villegas et al. (2023). This framework has introduced the notion of non-integer dimension tied to a constant entropy-loss rate across scales Poggialini et al. (2025). Crucially, this enables a distinction between scale-free and scale-invariant networks, with the latter admitting structural universality classes characterized by a well-defined spectral dimension, $d_{s}$. This global property of networks is related, for example, to the infrared singularity of the Gaussian process Cassi (1996), and provides a robust generalization of the standard concept of Euclidean dimension for heterogeneous systems Cassi (1992, 1996). Furthermore, structural scale invariance leads to the emergence of characteristic strange attractors in reciprocal space, allowing the identification of hidden symmetries and local dimensionalities undetectable hereto Villegas (2025). Here we exploit the LRG to tackle a crucial open question: can topological perturbations induce phase transitions in the very dimension of lattices and complex networks?
@@
Topological scale-invariant structures were defined in Poggialini et al. (2025), based on the LRG Villegas et al. (2023); Gabrielli et al. (2025), which analyzes the spectral properties of the network Laplacian $\hat{L}=\hat{D}-\hat{A}$, where $\hat{A}$ is the adjacency matrix and $\hat{D}$ is the diagonal degree matrix. Using the diffusion operator $e^{-\tau\hat{L}}$ as the evolution of a heat kernel at scale/diffusion time $\tau$, one defines the Laplacian density matrix De Domenico and Biamonte (2016) $\hat{\rho}(\tau)=\frac{e^{-\tau\hat{L}}}{Z(\tau)}$, with $Z(\tau)=\mathrm{Tr}\left[e^{-\tau\hat{L}}\right],$ playing the role of a partition function Villegas et al. (2022). This allows for a statistical mechanical description of the network structure, in which $\tau$ plays the role of inverse temperature and $\hat{L}$ acts as a Hamiltonian. The associated entropy $S(\tau)=-\mathrm{Tr}[\hat{\rho}(\tau)\log\hat{\rho}(\tau)]$ quantifies the effective number of active degrees of freedom at a given scale $\tau$, decreasing from $S=\log N$ at $\tau=0$ to $S=0$ at $\tau\rightarrow\infty$. Its derivative with respect to $\log\tau$ defines the entropic susceptibility or heat capacity, $C(\tau)=-\frac{dS}{d\log\tau}$, which measures the rate of information loss as the system is coarse-grained Villegas et al. (2022, 2023).
@@
Scale invariance is defined by the condition that $C(\tau)$ remains constant across a broad range of scales, indicating the absence of a characteristic structural scale in the system Poggialini et al. (2025). This is reflected by an entropy-loss rate proportional to the spectral dimension $d_{s}$, a global geometric quantity that generalizes the notion of Euclidean dimension to networks Cassi (1992, 1996). Hence, the presence of a plateau in the heat capacity, $C_{0}=d_{s}/2$, reveals the spectral dimension $d_{s}$ that characterizes the scale-invariant nature of homogeneous lattices or networks. Moreover, $C$ usually presents small-scale resonant modes at short times, a signature of the elementary small-scale structure of the lattice or network Falsi et al. (2024), i.e., the ultraviolet cut-off $\Lambda$ (in the jargon of the Renormalization Group). For the case of a 2D square lattice, this corresponds to the peak of $C$ around $\tau^{*}=1.24$ of the black dashed line in Fig. 1(a).
@@
The second peak at larger values of $\tau$ instead reflects the whole lattice scale. In particular, the recent categorization of structural fixed points has shed light on four main classes of scale-invariant networks Poggialini et al. (2025): (i) regular lattices, (ii) connected loopless networks, i.e., trees Donetti and Destri (2004), (iii) those with a non-vanishing clustering coefficient, as (u,v) flowers Rozenfeld et al. (2007) (including the Dorogovtsev-Goltsev-Mendes network Dorogovtsev et al. (2002) (DGM)), and Kim and Holme networks Holme and Kim (2002) and, (iv) hierarchical modular networks (HMNs) originally proposed taking inspiration from brain networks Moretti and Muñoz (2013).
@@
The ordered set of Laplacian eigenvalues provides further geometric insight, as it represents network frequency modes analogous to the usual Fourier analysis Villegas et al. (2023); Villegas (2025). Each eigenvalue represents a characteristic oscillation scale: low eigenvalues correspond to large-scale (low-frequency) modes, while high ones capture small-scale fluctuations (high frequencies). The spectral dimension inferred from entropy loss governs Laplacian embeddings, allowing any network to be unfolded into a low-dimensional manifold in $\kappa$-space. Using the first $n$ nontrivial Laplacian eigenvectors (for any connected network, the smallest eigenvalue, $\lambda_{0}=0$, is uniform and does not carry structural information), it is possible to obtain reliable $n$D representations, offering a stable embedding framework for analyzing network geometry. In these embeddings, scale-invariant networks display attractor-like structures where the correlation dimension $D$ matches the Euclidean dimension $d_{E}$ for lattices, and $D=d_{S}/2$ for heterogeneous networks Villegas (2025). This provides a minimal and robust geometric representation once considering the immediate embedding dimension after the correlation one Villegas (2025).
@@
Figure 1: Shortcuts in 2D square lattices. (a) Heat capacity, $C$, versus diffusion time, $\tau$, for different rewiring probabilities (see colorbar, $L=64$). The black dashed line corresponds to the unperturbed lattice. Inset: Zoom of the $\Lambda$-region. (b) First derivative of the heat capacity at the first peak versus rewiring probability, $p_{r}$, for different lattice sizes (see legend). (c) Finite-size scaling analysis of the vanishing peaks in (b). The distance of the size-dependent peak locations $p_{r,c}(N)$ from their asymptotic value for $N\rightarrow\infty$, $p_{r,c}^{\infty}$, scales as a power law of the system size, only at $p_{r,c}^{\infty}=0.10(1)$, revealing the existence of true scaling at criticality. (d) Low-dimensional representation using the three smallest low-frequency normalized network modes as the coordinate axes for a 2D square lattice of size $L=256$. Parameters: $p_{r}=2.3\times 10^{-4}$, $p_{r}=0$ (left inset), and $p_{r}=0.1$ (right inset). All curves have been averaged over $10^{3}-10^{5}$ realizations.
@@
We first consider the general case of regular lattices, where long-range connections are introduced by rewiring a fraction of the links. This is done by independently considering each endpoint of the links and randomly reconnecting it with a probability $p_{r}$, while preserving the total number of edges in the network. Fig. 1(a) shows the heat capacity as a function of the diffusion time for a 2D square network with varying rewiring probabilities. From now on, we consider the resulting giant component of the perturbed structure. As the rewiring probability increases, the plateau in the heat capacity narrows, until the early-time heat capacity peak vanishes, signaling the complete loss of translational symmetry (see the saddle point on the solid black curve in the inset). Considering this first peak as a hallmark of lattice-scale order, we quantify this transition by examining the derivative of the heat capacity at the characteristic time of the first peak (see Fig. 1(b)). Through finite-size scaling analyses, as reported in Fig. 1(c), we confirm the existence of a finite critical rewiring probability in the thermodynamic limit, $p_{r,c}^{sq}=0.10(1)$, which naturally defines the attraction basin of the 2D square lattice under the introduction of shortcuts. This geometric deformation can be further visualized using the 3D Laplacian projection introduced above. As reported in Fig. 1(d), the 2D lattice becomes distorted due to long-range connections, leading to a collapsed Gaussian-like network at the critical value (see insets of Fig. 1(d) and Supplementary Videos). A similar behavior is observed for other regular lattices, such as 2D triangular and hexagonal lattices (see Supplemental Material (SM) SM ), although with different critical rewiring thresholds: $p_{r,c}^{tr}=0.17(4)$ and $p_{r,c}^{hex}=0.055(10)$, respectively. Taken together, the disappearance of the short-scale peak in the heat capacity and the Gaussian collapse observed in Laplacian space constitute complementary signatures of a structural phase transition. These jointly define the limits of the attraction basin of the LRG structural fixed points.
@@
We now turn to the role of link dilution, a classical model for quenched disorder known to influence critical behavior in various universality classes Vojta (2006). Starting from a regular lattice, we remove links randomly with probability $p_{d}$. Fig. 2(a) displays the Laplacian embedding of a 2D squared lattice at different values of $p_{d}$ up to the percolation threshold, $p_{c}$, where the giant connected component disappears ($p_{c}=0.5$ for the 2D square lattice), and a random tree (RT) structure emerges. This is a direct consequence of the Alexander-Orbach conjecture Alexander, S. and Orbach, R. (1982), which predicts that, at criticality, the giant component for isotropic percolation has a universal spectral dimension $d_{s}=\nicefrac{{4}}{{3}}$ across all dimensions.
@@
To better characterize this phase transition, we also compute the correlation integral as defined by Grassberger and Procaccia Grassberger and Procaccia (1983); Villegas et al. (2024); Villegas (2025). Considering the projection of the $i^{th}$ node onto the $d$-dimensional space spanned by the first $d$ non-trivial Laplacian eigenvectors, $\mathbf{x}_{i}=\sqrt{N}\left(\langle\lambda_{1}|i\rangle,\langle\lambda_{2}|i\rangle,\ldots,\langle\lambda_{d}|i\rangle\right)$, we can define the correlation integral in the reciprocal space as $C(\ell)=\frac{2}{N(N-1)}\sum_{i>j}\Theta(\ell-||\mathbf{x_{i}-\mathbf{x_{j}}}||)$, where $\Theta$ is the Heaviside step function. If the scaling law $C(\ell)\propto\ell^{D}$ holds, then $D$ defines the network correlation dimension, which enables the detection of significant topological alterations in the perturbed network. As shown in Fig. 2(d), $D$ remains stable up to $p_{d,c}$, i.e., for low dilution, the lattice does not lose its bidimensional structure at large scale. From that point on, the dimensionality progressively decreases until it reaches the expected dimension for a RT.
@@
Figure 2: Diluted 2D square lattices. (a) Low-dimensional representation using the three smallest low-frequency normalized network modes as the coordinate axes, for a lattice with $p_{d}=0.35$, $p_{d}=0.2$ (left inset), and $p_{d}=0.5$ (right inset). (b) Heat capacity, $C$, versus diffusion time, $\tau$, for different dilute probabilities (see colorbar). Gray lines represent the expected plateau for a 2D lattice and a RT. (c) Second derivative of the heat capacity at the first peak versus dilute probability, $p_{d}$, for different lattice sizes (see legend). The vertical gray line represents the point where the 2D ultraviolet cut-off is expected to disappear. (d) Estimated correlation dimension, $D$, versus dilution probability, $p_{d}$ for different lattice sizes. Inset: Correlation integral $C(\ell)$ vs distance $\ell$ for different dilute probabilities. The gray area represents dilute probabilities between $p_{d,c}$ and the percolation threshold. All curves have been averaged over $10^{3}-10^{4}$ realizations.
@@
Both analyses capture the same phenomenon, yielding compatible estimates of the critical dilution probability $p_{d,c}=0.20(5)$ for the 2D square lattice. Again, an equivalent behavior emerges for triangular and hexagonal lattices (see SM SM ), but with different critical point, namely $p_{d,c}^{tr}=0.25(1)$ and $p_{d_{c}}^{hex}=0.11(2)$, respectively. For 3D cubic lattices, we have estimated $p_{r,c}=0.20(5)$, and $p_{d,c}=0.55(5)$ (see SM SM ). Note that performing correlation dimension analyses in the presence of shortcuts is considerably more challenging. While the method remains valid, the scaling region contracts, statistical noise increases, and extensive averaging becomes necessary. Moreover, the loss of spatial locality impairs the accuracy of Laplacian embeddings, making it more difficult to identify an effective embedding dimension. Conversely, in dilute models, it is another complementary measure that allows for the analysis of increasing system sizes, as it depends only on a few eigenvectors.
@@
We now extend our analysis to heterogeneous structures. In the case of generic trees, the critical rewiring threshold scales as $p_{r,c}\sim\mathcal{O}(\nicefrac{{1}}{{N}})$, while link deletion leads to a trivial threshold $p_{d,c}=1$, since the spectral dimension remains unchanged. In contrast, DGM networks exhibit more significant and non-trivial behavior. For these, we identify critical rewiring and dilution thresholds $p_{r,c}\neq 0$ (see SM SM for a discussion on the issue) and $p_{d,c}=0.15(5)$, respectively (the percolation threshold is $p_{c}=0.95(5)$, see SM SM ). Figs. 3(a) and (c) show the correlation dimension, $D$, as a function of $p_{d}$ for DGM and HMN networks. In DGM networks, $D$ remains constant up to $p_{d,c}$ and then begins to decrease. In particular, these networks show evidence of flowing to the vicinity of an unstable structural fixed point for $p_{d}=0.75(5)$: as shown in Fig. 3(b), their structure turns out to be analogous to a BA network, following a power-law degree distribution $P(\kappa)\sim\kappa^{-3}$. Furthermore, the heat capacity stabilizes at $C_{0}=1$, matching the BA spectral signature Villegas et al. (2023); Poggialini et al. (2025) (see Fig. 3(b)). In HMNs, a similar decay in $D$ is observed beyond the corresponding $p_{d,c}$. However, unlike DGM networks, after the percolation threshold, $D$ stabilizes at the RT spectral dimension, $d_{s}=4/3$. The absence of hubs prevents the system from flowing close to the BA unstable fixed point reported in the DGM case.
@@
Figure 3: Diluted scale-invariant networks. Estimated correlation dimension, D, versus dilution probability, $p_{d}$, for network of different levels, s (see legend), in the case of: (a) DGM ($N_{s}=\frac{3+3^{s}}{2}$) and (c) HMN with $m_{0}=3$ and $\alpha=4$ ($N_{s}=2^{s}m_{0}$). (b) Degree distribution versus node connectivity, $\kappa$, for a DGM network with $p_{d}=0$(blue points) and $p_{d}=0.75$. The black dashed line reports the solution for a BA network, $P(\kappa)\propto\kappa^{-3}$. Inset: Specific heat for different dilution probabilities. For $p_{d}=0.75$, the network exhibits a plateau on $C_{0}=1$ as expected for BA networks. (d) Lacunarity index for different diluted networks and lattices for a fixed scale $\epsilon=10^{-2}$. All curves have been averaged over $10^{3}-10^{4}$ realizations.
@@
Finally, to further characterize the loss of structural regularity, we compute the lacunarity Mandelbrot (1983); Gabrielli et al. (2006), $\lambda_{\varepsilon}$, or void distribution, defined as the squared coefficient of variation of pixels per box in a box-counting analysis: $\lambda_{\varepsilon}=\frac{\sigma_{\varepsilon}^{2}}{\mu_{\varepsilon}^{2}}$, with $\varepsilon$ the box-size. As seen in Fig. 3(d), by using a 3D embedding space, lacunarity increases with dilution across all studied networks. This indicates a breakdown of statistical self-similarity and the emergence of heterogeneity in the spatial complexity. For HMNs, high lacunarity is present even without dilution, indicating their intrinsic geometric irregularity. This is particularly significant because when lacunarity is low, fractals behave similarly to analytically continued hypercubic lattices Gefen et al. (1983). Therefore, lacunarity may serve as a signature of anomalous universal behavior.

### v2

@@ Abstract
Dimension in physical systems determines universal properties at criticality. Yet, the impact of structural perturbations on dimensionality remains largely unexplored. Here, we characterize the attraction basins of structural fixed points in scale-invariant networks from a renormalization group perspective, demonstrating that basin stability connects to a structural phase transition. This topology-dependent effect, which we term geometric criticality, triggers a geometric breakdown hitherto unknown, which induces non-trivial fractal dimensions and unveils hidden LRG flows toward unstable structural fixed points. Our systematic study of how networks and lattices respond to disorder paves the way for future analysis of non-ergodic behavior induced by quenched disorder.
@@
Topological shortcuts and structural sparsity shape complex dynamics across a broad spectrum of physical systems [1, 2, 3]. Shortcuts are key, for instance, in shaping brain functionality [4] or in improving the speed, efficiency, robustness, and functional diversity across neural networks [5, 6, 7]. Similar concepts pervade other areas, such as wormholes in general relativity [8] or wormhole attacks in wireless sensor networks [9]. In contrast, sparsity – for example, selective bond breaking in zeolites [10, 11] – limits connectivity and tends to suppress large-scale organization. Vacancies – missing links in real materials – often compromise emergent phases by disrupting connectivity and collective behavior [3]. Such defects are known to modify electronic, magnetic, and structural properties in oxides, alloys, and skyrmion crystals [12, 13, 14]. Despite their relevance, understanding whether shortcuts and sparsity lead to structural transitions in complex systems remains a crucial open challenge.
@@
In the context of complex networks, shortcuts were first formalized in the celebrated Watts-Strogatz (WS) model [15]. There, a simple ring lattice is ‘rewired’ to introduce increasing amounts of disorder, allowing for a middle ground between regular and random networks while preserving high clustering and a very slowly growing diameter [16, 17]. Extensions of the WS model, where the rewiring probability decays with distance, have further demonstrated that such shortcuts can modify the spectral dimension of the underlying graph [18]. However, do small-world effects merely induce smooth crossovers towards randomness or possibly lead to genuine structural phase transitions?
@@
Dilute lattice models have long suggested that structural perturbations may influence universality classes [19]. Griffiths theoretically demonstrated that disordered systems could exhibit non-analytic free energy over an extended temperature range [20]. Understanding Griffiths phases is essential for comprehending many other disordered systems, such as dilute ferroelectric materials [21, 22]. However, no general framework has emerged to predict when and how such quenched disorder induces non-ergodic behavior in arbitrary network architectures.
@@
This motivated recent work on intrinsically scale-invariant networks, where structural fixed points were defined through the Laplacian Renormalization Group (LRG) [23]. This framework has introduced the notion of non-integer dimension tied to a constant entropy-loss rate across scales [24]. Crucially, this enables a distinction between scale-free and scale-invariant networks, with the latter admitting structural universality classes characterized by a well-defined spectral dimension, $d_{s}$. This global property of networks is related, for example, to the infrared singularity of the Gaussian process [25], and provides a robust generalization of the standard concept of Euclidean dimension for heterogeneous systems [26, 25]. Furthermore, structural scale invariance leads to the emergence of characteristic strange attractors in reciprocal space, allowing the identification of hidden symmetries and local dimensionalities undetectable via standard methods [27]. Here, we exploit the LRG to address a fundamental question: Is it possible to define attraction basins for lattices and scale-invariant networks by exploring their response to topological perturbations?
@@
Topological scale-invariant structures were defined in [24], based on the LRG [23, 28], which analyzes the spectral properties of the network Laplacian $\hat{L}=\hat{D}-\hat{A}$, where $\hat{A}$ is the adjacency matrix and $\hat{D}$ is the diagonal matrix of node degrees. Using the diffusion operator $e^{-\tau\hat{L}}$ as the evolution of a heat kernel at scale/diffusion time $\tau$, one defines the Laplacian density matrix [29] $\hat{\rho}(\tau)=\frac{e^{-\tau\hat{L}}}{Z(\tau)}$, with $Z(\tau)=\mathrm{Tr}\left[e^{-\tau\hat{L}}\right$], playing the role of a partition function [30]. This allows for a statistical mechanical description of the network structure, in which $\tau$ plays the role of inverse temperature and $\hat{L}$ acts as a Hamiltonian. The associated entropy $S(\tau)=-\mathrm{Tr}[\hat{\rho}(\tau)\log\hat{\rho}(\tau)]$ quantifies the effective number of active degrees of freedom at a given scale $\tau$, decreasing from $S=\log N$ at $\tau=0$ to $S=0$ at $\tau\rightarrow\infty$. Its derivative with respect to $\log\tau$ defines the entropic susceptibility or heat capacity, $C(\tau)=-\frac{dS}{d\log\tau}$, which measures the rate of information loss as the system is coarse-grained [30, 23].
@@
Scale invariance is defined by the condition that $C(\tau)$ remains constant across a broad range of scales, indicating the absence of a characteristic structural scale in the system [24]. This implies an entropy-loss rate proportional to the spectral dimension $d_{s}$, a global geometric quantity that generalizes the notion of Euclidean dimension to networks [26, 25]. Hence, the presence of a plateau in the heat capacity, $C_{0}=d_{s}/2$, reveals the spectral dimension $d_{s}$ that characterizes the scale-invariant nature of homogeneous lattices or networks. Moreover, $C$ usually presents small-scale resonant modes at short times, a signature of the elementary small-scale structure of the lattice or network [31], i.e., the ultraviolet cut-off $\Lambda$ (in the jargon of the Renormalization Group). For a 2D square lattice, this corresponds to the peak of $C$ at $\tau\approx 1.24$ (black dashed line in Fig. 1(a)).
@@
The second peak at larger values of $\tau$ instead reflects the whole lattice scale. In particular, the recent categorization of structural fixed points has shed light on four main classes of scale-invariant networks [24]: (i) regular lattices, (ii) connected loopless networks, i.e., trees [32], (iii) those with a non-vanishing clustering coefficient, as (u,v) flowers [33] (including the Dorogovtsev-Goltsev-Mendes network [34] (DGM)), and Kim and Holme networks [35] and, (iv) hierarchical modular networks (HMNs) originally proposed taking inspiration from brain networks [36].
@@
The ordered set of Laplacian eigenvalues provides further geometric insight, as it represents network frequency modes analogous to the usual Fourier analysis [23, 27]. Each eigenvalue represents a characteristic oscillation scale: low eigenvalues correspond to large-scale (low-frequency) modes, while high ones capture small-scale fluctuations (high frequencies). The spectral dimension inferred from entropy loss governs Laplacian embeddings, allowing any network to be unfolded into a low-dimensional manifold in $\kappa$-space. Using the first $n$ nontrivial Laplacian eigenvectors (for any connected network, the smallest eigenvalue, $\lambda_{0}=0$, is uniform and does not carry structural information), it is possible to map network nodes into $\mathbb{R}^{n}$, providing a geometric representation. In these embeddings, scale-invariant networks display attractor-like structures where the correlation dimension $D$ matches the Euclidean dimension $d_{E}$ for lattices, and $D=d_{S}/2$ for heterogeneous networks [27]. This provides a minimal and robust geometric representation just considering the minimum integer embedding dimension $d_{E}\geq D$ [27].
@@
Figure 1: Shortcuts in 2D square lattices. (a) Heat capacity, $C$, versus diffusion time, $\tau$, for different rewiring probabilities (see colorbar, $L=64$). The black dashed line corresponds to the unperturbed lattice. Inset: Zoom of the $\Lambda$-region. (b) First derivative of the heat capacity at the first peak versus rewiring probability, $p_{r}$, for different lattice sizes (see legend). (c) Finite-size scaling analysis of the vanishing peaks in (b). The distance of the size-dependent peak locations $p_{r,c}(N)$ from their asymptotic value for $N\rightarrow\infty$, $p_{r,c}^{\infty}$, scales as a power law of the system size, only at $p_{r,c}^{\infty}=0.10(1)$, revealing the existence of true scaling at criticality. (d) Low-dimensional representation using the three smallest low-frequency normalized network modes as the coordinate axes for a 2D square lattice of size $L=256$. Parameters: $p_{r}=2.3\times 10^{-4}$, $p_{r}=0$ (left inset), and $p_{r}=0.1$ (right inset). All curves have been averaged over $10^{3}$$10^{5}$ realizations.
@@
We first consider the general case of regular lattices, where long-range connections are introduced by rewiring a fraction of the links. This is achieved by independently rewiring each link endpoint with probability $p_{r}$, while preserving the total number of edges in the network. Fig. 1(a) shows the heat capacity as a function of the diffusion time for a 2D square network with varying rewiring probabilities. From now on, we consider the resulting giant component of the perturbed structure. As the rewiring probability increases, the plateau in the heat capacity narrows, until the early-time heat capacity peak vanishes, signaling the breakdown of translational symmetry (see the saddle point on the solid black curve in the inset). Considering this first peak as a hallmark of lattice-scale order, we quantify this transition by examining the derivative of the heat capacity at the characteristic time of the first peak (see Fig. 1(b)). Through finite-size scaling analyses, as reported in Fig. 1(c), we confirm the existence of a finite critical rewiring probability in the thermodynamic limit, $p_{r,c}^{sq}=0.10(1)$, which naturally defines the attraction basin of the 2D square lattice under the introduction of shortcuts. This geometric deformation can be further visualized using the 3D Laplacian projection introduced above. As reported in Fig. 1(d), the 2D lattice becomes distorted due to long-range connections, leading to a collapsed Gaussian-like network at the critical value (see insets of Fig. 1(d) and Supplementary Videos [37]). A similar behavior is observed for other regular lattices, such as 2D triangular and hexagonal lattices (see Appendix B), although with different critical rewiring thresholds: $p_{r,c}^{tr}=0.17(4)$ and $p_{r,c}^{hex}=0.055(10)$, respectively. Taken together, the disappearance of the short-scale peak in the heat capacity and the Gaussian collapse observed in Laplacian space constitute complementary signatures of a structural phase transition. These jointly define the limits of the attraction basin of the LRG structural fixed points.
@@
We now turn to the role of link dilution, a classical model for quenched disorder known to influence critical behavior in various universality classes [19]. Starting from a regular lattice, we remove links randomly with probability $p_{d}$. Fig. 2(a) displays the Laplacian embedding of a 2D square lattice at different values of $p_{d}$ up to the percolation threshold, $p_{c}$, where the giant connected component disappears ($p_{c}=0.5$ for the 2D square lattice), and a random tree (RT) structure emerges. This is a direct consequence of the Alexander-Orbach conjecture [38], which predicts that, at criticality, the giant component for isotropic percolation has a universal spectral dimension $d_{s}=\nicefrac{{4}}{{3}}$ across all dimensions.
@@
To better characterize this phase transition, we also compute the correlation integral as defined by Grassberger and Procaccia [39, 40, 27]. Using the projection of the ithi-{th} node onto the $d$-dimensional space spanned by the first $d$ non-trivial Laplacian eigenvectors, $\mathbf{x}_{i}=\sqrt{N}\left(\langle\lambda_{1}|i\rangle,\langle\lambda_{2}|i\rangle,\ldots,\langle\lambda_{d}|i\rangle\right)$, we can define the correlation integral in the reciprocal space as $C(\ell)=\frac{2}{N(N-1)}\sum_{i>j}\Theta(\ell-\|\mathbf{x}_{i}-\mathbf{x}_{j}\|)\sim\ell^{D}$, where $\Theta$ is the Heaviside step function. If the scaling law $C(\ell)\propto\ell^{D}$ holds, then $D$ defines the network correlation dimension, which enables the detection of significant topological alterations in the perturbed network. As shown in Fig. 2(d), $D$ remains stable up to $p_{d,c}$, i.e., for low dilution, the lattice does not lose its bidimensional structure at large scale. From that point on, the dimensionality progressively decreases until it reaches the expected dimension for a RT.
@@
Figure 2: Diluted 2D square lattices. (a) Low-dimensional representation using the three smallest low-frequency normalized network modes as the coordinate axes, for a lattice with $p_{d}=0.35$, $p_{d}=0.2$ (left inset), and $p_{d}=0.5$ (right inset). (b) Heat capacity, $C$, versus diffusion time, $\tau$, for different dilute probabilities (see colorbar). Gray lines represent the expected plateau for a 2D lattice and a RT. (c) Second derivative of the heat capacity at the first peak versus dilute probability, $p_{d}$, for different lattice sizes (see legend). The vertical gray line represents the point where the 2D ultraviolet cut-off is expected to disappear. (d) Estimated correlation dimension, $D$, versus dilution probability, $p_{d}$ for different lattice sizes. Inset: Correlation integral $C(\ell)$ vs distance $\ell$ for different dilute probabilities. The gray area represents dilute probabilities between $p_{d,c}$ and the percolation threshold. All curves have been averaged over $10^{3}$$10^{4}$ realizations.
@@
Both analyses capture the same phenomenon, yielding compatible estimates for the critical dilution probability $p_{d,c}=0.20(5)$ for the 2D square lattice. Again, an equivalent behavior emerges for triangular and hexagonal lattices (see Appendix B), but with different critical points: $p_{d,c}^{tr}=0.25(1)$ and $p_{d_{c}}^{hex}=0.11(2)$, respectively. For 3D cubic lattices, we estimate a critical dilution threshold of $p_{d,c}=0.55(5)$, while the rewiring threshold was found to be $p_{r,c}=0.20(5)$ (see Appendix C). Note that performing correlation dimension analyses in the presence of shortcuts is considerably more challenging. While the method remains valid, the scaling region contracts, statistical noise increases, and extensive averaging becomes necessary. Moreover, the loss of spatial locality impairs the accuracy of Laplacian embeddings, making it more difficult to identify an effective embedding dimension. Conversely, in dilute models, it is another complementary measure that allows for the analysis of increasing system sizes, as it depends only on a few eigenvectors.
@@
We now extend our analysis to heterogeneous structures. For generic trees, the critical rewiring threshold scales as $p_{r,c}\sim\mathcal{O}(\nicefrac{{1}}{{N}})$, while link deletion leads to a trivial threshold $p_{d,c}=1$, since the spectral dimension remains unchanged (see Appendix E). In contrast, DGM networks exhibit more significant and non-trivial behavior. For these, we identify non-trivial critical thresholds for rewiring and dilution: $p_{r,c}\neq 0$ (see Appendix D for a discussion on the issue) and $p_{d,c}=0.15(5)$, respectively (the percolation threshold is $p_{c}=0.95(5)$, see Appendix D). Figs. 3(a) and (c) show the correlation dimension, $D$, as a function of $p_{d}$ for DGM and HMN networks. In DGM networks, $D$ remains constant up to $p_{d,c}$ and then begins to decrease. In particular, these networks show evidence of flowing to the vicinity of an unstable structural fixed point for $p_{d}=0.75(5)$: as shown in Fig. 3(b), their structure converges to that of a Barabási–Albert (BA) network, following a power-law degree distribution $P(\kappa)\sim\kappa^{-3}$. Furthermore, the heat capacity stabilizes at $C_{0}=1$, matching the BA spectral signature [23, 24] (see Fig. 3(b)). In HMNs (see Appendix F for further analysis), a similar decay in $D$ is observed beyond the corresponding $p_{d,c}$. However, unlike DGM networks, after the percolation threshold, $D$ stabilizes at the RT spectral dimension, $d_{s}=4/3$. The absence of hubs prevents the system from flowing close to the BA unstable fixed point reported in the DGM case.
@@
Figure 3: Diluted scale-invariant networks. (a, c) Estimated correlation dimension, $D$, versus dilution probability, $p_{d}$, for networks of different generation levels, $s$ (see legend), for: (a) DGM ($N_{s}=\frac{3+3^{s}}{2}$) and (c) HMN with $m_{0}=3$ and $\alpha=4$ ($N_{s}=2^{s}m_{0}$). (b) Degree distribution, $P(\kappa)$, versus node degree, $\kappa$, for a DGM network with $p_{d}=0$ (blue points) and $p_{d}=0.75$ (orange points). The black dashed line indicates the power-law scaling of a BA network, $P(\kappa)\propto\kappa^{-3}$. Inset: Specific heat for different dilution probabilities; for $p_{d}=0.75$, the plateau at $C_{0}=1$ confirms the BA limit. (d) Lacunarity index, $\lambda$, calculated at a fixed scale $\epsilon=10^{-2}$ for different network topologies (x-axis). Colors distinguish different dilution probabilities (see legend). The horizontal dashed line marks the theoretical reference $\lambda=1$ (where variance equals mean). All curves have been averaged over $10^{3}$$10^{4}$ realizations.
@@
Finally, to further characterize the loss of structural regularity, we compute the lacunarity [41, 42], $\lambda_{\varepsilon}$, or void distribution, defined as the square coefficient of variation of pixels per box in a box-counting analysis: $\lambda_{\varepsilon}=\frac{\sigma_{\varepsilon}^{2}}{\mu_{\varepsilon}^{2}}$, with $\varepsilon$ the box-size. As seen in Fig. 3(d), by using a 3D embedding space, lacunarity increases with dilution across all studied networks. This indicates a breakdown of statistical self-similarity and the emergence of heterogeneity in the spatial complexity. For HMNs, high lacunarity is present even without dilution, indicating their intrinsic geometric irregularity. This is particularly significant because when lacunarity is low, fractals behave similarly to analytically continued hypercubic lattices [43]. Therefore, lacunarity may serve as a signature of anomalous universal behavior.